a bicycle chain moves along two sprockets when you shift gears. the number of teeth on the front sprocket form a ratio. equivalent ratios provide equal pedaling power. find a ratio equivalent to the ratio 52 over24

Answer by richwmiller(17219)

52/24=26/12=13/6